OpenStudy (anonymous):

Match each expression with its solution. Match Term Definition seven point five times one hundred A) seven thousand five hundred zero point seven five times ten B) seven hundred fifty seventy five thousand divided by ten C) seven point five seven thousand five hundred divided by one hundred D) seventy-five

OpenStudy (anonymous):

@Chineseboy15

OpenStudy (anonymous):

7.5 *100 750

OpenStudy (anonymous):

First change all these numbers in to Arabic numbers.

OpenStudy (anonymous):

You do this or I do this?

OpenStudy (anonymous):

7.5 * 100 a) 7500 0.75 * 10 B) 750 75000 / 10 C) 7.5 7500 / 100 D) 75

OpenStudy (anonymous):

is that what you wanted??

OpenStudy (anonymous):

Great.

OpenStudy (anonymous):

7.5*100=750 0.75*10=7.5 75000/10=7500 7500/100=75

OpenStudy (anonymous):

so the answers are A B C D

OpenStudy (anonymous):

well 1 = A 2 = B 3 = C 4 = D

OpenStudy (anonymous):

The answer is B C A D. Sorry.

OpenStudy (anonymous):

7.5*100=750 B) B) 750 0.75*10=7.5 C) C) 7.5 75000/10=7500 A) a) 7500 7500/100=75 D) D) 75